OpenStudy (anonymous):

Graph the ellipse and locate the foci. x^2/36 + y^2/11 = 1 help please?

OpenStudy (anonymous):

I did one of these just last week but now I'm lost again

OpenStudy (anonymous):

@beccaboo333 do you know how to solve? If you do I'd appreciate a walkthrough

OpenStudy (beccaboo333):

If I was good at math, I'd help you.. But I'm not.. Uhhh, @tHe_FiZiCx99 @ganeshie8 @ikram002p

OpenStudy (mathmate):

 Standard formula for ellipse centred at origin: \(\large \frac{x^2}{a^2}+\frac{y^2}{b^2}=1\) c=distance from centre to focus = \(sqrt(|b^2-a^2|)\) Axis (line joining foci) is horizontal when a>b Axis is vertical when b>a Compare standard equation to find a and b (distances from centre to vertices)